How many grams of carbon dioxide would be formed if 38.9 g C2H2 reacted completely with oxygen?
1
Expert's answer
2020-10-16T08:26:40-0400

Given 38.9g of C2H2 completely reacted with O2

equation: 2C2H2 + 5O2 = 4CO2 + 2H2O

moles of C2H2 = 38.9/26.0 = 1.496

mole ratio of C2H2 to CO2 = 2:4

moles of CO2 = 4*1.496/2 = 2.992 moles

mass of CO2 produced = 2.992 moles * 44 g/mol= 131.662g